OpenStudy (anonymous):

if y varies inversely as x and x=50 when y=2.7, find y when x=80.( round off your answer to the nearest tenth)

OpenStudy (anonymous):

y = k*(1/x) 2.7 = k*(1/50) solve for k 2.7/(1/50) = k k = 135 no use your value for k to solve for y y = 135*(1/80) y = 135/80 y = 1.6875 round to nearest 10th y = 1.7